Presented below are two independent situations.

1. On January 1, 2020, Tamarisk Company issued $312,000 of 9%, 10-year bonds at par. Interest is payable quarterly on April 1, July 1, October 1, and January 1.
2. On June 1, 2020, Vaughn Company issued $264,000 of 11%, 10-year bonds dated January 1 at par plus accrued interest. Interest is payable semiannually on July 1 and January 1.


For each of these two independent situations, prepare journal entries to record the following. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

(a) The issuance of the bonds.
(b) The payment of interest on July 1.
(c) The accrual of interest on December 31.
Answer #1
Answer
1
Journal
Date Account title Debit Credit
a) Jan 1. 2020 Cash $     312,000
Bonds payable $     312,000
b) July 1. Bond interest expense $         7,020
Cash $         7,020
c) Dec 31. Bond interest expense $         7,020
Bond interest payable $         7,020
Quarterly interest payment = 312,000 x 9/100 x 3/12 = $ 7,020
2.  
Journal
Date Account title Debit Credit
a) June 1. 2020 Cash $     265,200
Bonds payable $     264,000
Interest payable $         1,200
b) July 1. Bond interest expense $       14,520
Cash $       14,520
c) Dec-31 Bond interest expense $       14,520
Bond interest payable $       14,520
Semi-annually interest payment = 264,000 x 11/100 x 6/12 = $14,520
Interest payable on June 1 = 264,000 x 11/100 x 5/12 = $12,100